# references.bib

### v1

@@ 1 Introduction
The square-root law for price impact is arguably one of the most robust empirical regularities discovered in the last 30 years – see [loeb1983trading, grinold2000active, almgren2005direct, toth2011anomalous, donier2015million, toth2016square] and [bouchaud2018trades, webster2023handbook] for reviews. It states that when executing a buy (resp. sell) meta-order of total size $Q$, sliced and diced into $N$ child orders of size $q=Q/N$, the price on average moves up (down) by an amount proportional to $\sqrt{Q}$. Price impact is, quite remarkably, found to be approximately independent of both $N$ and of the total time $T$ needed to achieve full execution [bucci_impact_vol2019]. In other words, provided the participation rate is not too large, average price impact only depends – to a first approximation – on the total volume traded $Q$, but not on execution schedule [bouchaud2018trades].
@@
Such a square-root dependence, and its apparent universality across a wide variety of markets [bouchaud2018trades], is surprising and non-intuitive. The classical Kyle model would rather predict a linear dependence of impact on $Q$, with a slope usually called “Kyle’s $\lambda$[kyle1985continuous]. Square-root impact can be interpreted as an effective $Q$-dependent $\lambda$ that diverges as $Q^{-1/2}$ when $Q\to 0$, as if there was a dearth of liquidity for small volumes [donier2016walras].
@@
Several theoretical ideas have been put forth in the literature to explain non-linear impact. Some models predict a concave price impact $Q^{\delta}$ with $\delta\leq 1$ related to the power-law tail exponent $\alpha$ of the executed volume [gabaix2006] or the power-law tail exponent $\gamma$ of the time autocorrelation of the sign of market orders [farmer2013efficiency, lillo2005theory]. However, as recently shown by Sato and Kanazawa [kanazawasato2024] using ID-resolved data from the Tokyo Stock Exchange (TSE), the predicted relations between $\delta$ and $\alpha$ or $\gamma$ are not borne out by the data: whereas $\alpha$ and $\gamma$ significantly differ between stocks, exponent $\delta$ remains stubbornly anchored around $\delta=1/2$, i.e. the value corresponding to the square-root law. Another possibility is to invoke the “propagator model” which describes the power-law decay of the impact of individual orders [bouchaud2003fluctuations, bouchaud2009markets, bouchaud2018trades]. However, in this case all the concavity comes from a time effect, not from a volume effect – a possibility clearly ruled out, see [bucci_impact_vol2019] and Fig. 6 below.
@@
Yet another line of thought builds upon the classical Glosten-Milgrom model [glosten1985bid] and assumes that metaorders are issued by traders possessing information about future prices, as is traditional in the economics literature [hasbrouck2007empirical, bouchaud2010impact]. Market makers try to detect these informed traders, but do not know their actual participation rate in the total order flow. Using Bayesian arguments, Saddier & Marsili [saddier2024bayesian] then derive a square-root impact law under certain (weak) conditions and predict that the impact will decay after the end of execution as $t^{-1/2}$ when $t\gg T$.
@@
This last prediction is shared with the Locally Linear Order Book (LLOB) model, which is based on a dynamical theory of “latent” liquidity, i.e. volumes that are intended to be exchanged but are not necessarily visible in the order book [donier2015fullyconsistentminimalmodel, bouchaud2018trades]. Plausible assumptions about the dynamics of such liquidity lead to a generic “V-shape” in the vicinity of the current price, meaning that local liquidity is vanishingly small – in line with our above comment on the divergence of Kyle’s $\lambda$ for small $Q$’s. Incoming buy (sell) trades then face more and more “resistance” as the price moves up (down). The linear increase of liquidity then translates into a square-root behaviour for price impact [donier2015fullyconsistentminimalmodel, donier2016walras]. However, a more precise formulation of the model suggests the existence of two regimes: a small $Q$ regime where the impact is linear and a larger $Q$ regime where the impact is square root. Whereas such a crossover between the two regimes has indeed been reported in Ref. [bucci2019crossover] for very small $Q$’s, the square-root regime sets in for volumes 300 times smaller than predicted by the theory. We will discuss this conundrum more in details below, in the light of our empirical results.
@@
Let us finally mention that the square-root impact law can also be justified using dimensional arguments coupled with additional plausible assumptions, see [pohl2017amazing, kyle2018market].
@@
Despite its critical importance for both financial microstructure and an asset pricing [gabaix2021search, bouchaud2022inelastic], the very origin of this central phenomenon remains a topic of debate. The universality of the phenomenon suggests a purely mechanical, rather than informational, origin, however this point is controversial and at odds with most of the economics literature on the subject, starting with the famous Kyle model [kyle1985continuous]. The aim of the present paper is to give more credence to the “mechanical hypothesis” of the square-root market impact law using an ID-resolved data set from the TSE. Our main results are the following:
@@
  • “Fast” traders, for which the holding period is less than a day, represent between 50% and 60% of the executed market orders. This includes market makers (HFT) and short term traders. Correspondingly, the fraction of market orders executed against “fast” traders represents nearly half of the exchanged volume, a number far too low to vindicate the standard interpretation of the square-root impact within the LLOB framework [benzaquen2017marketimpactmultitimescaleliquidity]. Since this empirical fact is inconsistent with the standard LLOB framework, it motivates us to propose a new interpretation of the LLOB framework, as presented in this paper.
  • One can define refill sequences for liquidity providers. One finds that the size of those sequences is also power-law distributed, as was found for metaorders in [sato2023inferring], following the suggestion of Lillo, Mike & Farmer [lillo2005theory].
@@ 2.1 A unique dataset
Our study is based on a dataset from the Tokyo Stock Exchange (TSE), provided by the Japan Exchange Group (JPX) for academic purposes only and already used in [sato2023inferring]. The dataset contains all orders sent to the exchange, with a unique order ID, a virtual server ID, the price and type of the order, the volume and price of the best quotes, for all stocks available on the exchange from 2012 to 2018. Here the virtual server ID is the unit of trading accounts on the TSE. Technically it is not a membership ID (i.e., the corporate level ID) because any trader may have several virtual servers to avoid the submission-number limit during a fixed interval. However, one can reconstruct an effective trader ID, called the Trading Desk, by properly aggregating those virtual server IDs (see [sato2023inferring, goshima2019trader] for the details). In this paper, the Trading Desks are referred to as trader IDs.
@@
Item (i) is particularly important regarding the claim that impact is an universal mechanism, independent of the type of traders and the information content of the trades [toth2017short, bouchaud2018trades]. Indeed, access to metaorders data is rare (the Ancerno dataset being one exception [zarinelli2015beyond, bucci2018slow, bucci2019crossover]), and most of datasets used in the literature are proprietary and may be plagued by conditioning effects [bouchaud2018trades, webster2023handbook]. [^1] Note however that CFM data [toth2011anomalous] was acquired in such a way to minimize conditioning effects such as those discussed in [bouchaud2018trades], ch. 12.3, see also [brokmann2015slow].
Item (ii) represents an interesting opportunity to categorize traders as market makers (MM), high frequency traders (HFT) or low frequency traders, see 2.3 and understand better their typical impact on the market. For example, the long term debate about the benefits of HFT for market stability has been dramatically improved with this kind of dataset [hosaka2014analysis, goshima2019trader]. In addition, these identifiers enabled us to analyze the behavior of liquidity providers, which is only possible if we have access to the ownership of all limit or cancellation orders, see Section 5.
@@ 2.2 General stylized facts about metaorders execution
We show in Fig. 1 (bottom graph) the distribution of executed volume fraction $f:=Q/V_{D}$, which shows a broad maximum in the region $f\in[0.1\%,1\%]$. Some metaorders correspond to $10\%$ of the daily volume but they are relatively rare. Similarly, very small metaorders of size $<0.01\%$ have a very small probability. The range $f\in[0.1\%,10\%]$ for metaorders is typical of firms like, e.g., AQR or CFM [AQR, brokmann2015slow].
The average time between child orders ${\Delta t}$ and the average number of child orders are plotted in Fig. 1 (top graph) as a function of $f=Q/V_{D}$, where $V_{D}$ is the executed volume during the day in shares.. We also show as a shaded region the standard deviation of these quantities. One sees that the average time between child orders mildly increases as a function of $f$, ranging from $25$ secs. for $f=0.01\%$ to $150$ secs. for $f=1\%$, before saturating or even slightly decreasing for larger values of $f$. Hence the average execution time $T$ increases slightly faster than $Q$ itself, except perhaps for the largest metaorders.
The increase of $\Delta t$ when $Q$ increases is related to the fact that child orders are more and more aggressive in order to complete execution, so traders wisely wait longer before sending the next one, lest they are detected by market makers. This however becomes difficult for large $f$s, because traders are also attempting to execute their metaorders as quickly as possible. Note that translated into total execution time $T$, these results show that for $f=0.01\%$, the typical value of $T$ is $\approx 200$ seconds, whereas for large metaorders with $f=10\%$, $T\approx$ 90 minutes. These numbers are however only indicative and the total duration of execution can rise to a full day.
The typical number $N$ of child orders is around $10$ for $f\lesssim 1\%$ before increasing steeply for larger $f$. This reflects the fact that the available volume at the best quotes is relatively small and if traders want to avoid “eating into the book”, then the size $q$ of child orders is also limited, which mechanically pushes the number of child orders up when $Q$ increases. Note that typically the volume available at the best quote is around $10^{-4}V_{D}$ for the most liquid stocks.
@@ 2.3 Time scales & Market ecology
These time scales are relevant to understand order flow and liquidity dynamics, and therefore price impact. Indeed, whereas market makers allow orderly trading by acting as intermediaries between final buyers and final sellers, their inventory constraint prevent them from offering “resistance” to large buy or sell metaorders. Only slow liquidity can counter-act persistent order imbalance [benzaquen2017marketimpactmultitimescaleliquidity].
@@
Using the reversal time $\tau$ as a criterion to separate “fast” and “slow” traders, we compute the contribution of the two categories to the trading activity using either the fraction of the market order volume executed by fast traders, $V_{\text{fast}}/V_{D}$ or the relative fraction of the number of fast traders, $N_{\text{fast}}/N_{D}$. $N_{\text{fast}}$ is the number of “fast” traders in a given session, whereas $N_{D}$ is the total number of traders having traded at least once during the session. We show in Fig. 3 the histogram of these ratios, computed over all sessions and all stocks of our database. Whereas the most probable value of $N_{\text{fast}}/N_{D}$ is around $8\%$, the most probable value of the ratio $V_{\text{fast}}/V_{D}$ is between $50\%$ and $60\%$. Therefore, the contribution of “slow” market orders to total volume is roughly one-half. This finding is in line with [Ohyama]. One can also determine the fraction of market orders executed against fast traders, which is found to be in the range $60\%$ to $70\%$.
@@
The conclusion of this study is that while fast traders are dominant in terms of volume, the contribution of slow volumes to trading activity actually of the same order of magnitude. This finding is important since the standard interpretation of the square-root impact law in terms of latent liquidity [donier2015fullyconsistentminimalmodel, benzaquen2017marketimpactmultitimescaleliquidity] assumes that slow volumes are a factor $\sim 300$ times smaller than $V_{D}$ [bucci2019crossover], which is certainly not the case here. In the next Section, we will revisit the empirical evidence for square-root impact and propose a new interpretation of the LLOB model.
@@ 3 Square-root impact: micro-scales & meso-scales
As recalled in the introduction, there is overwhelming empirical evidence for the square-root impact law for metaorders, which has again been confirmed in great detail in [kanazawasato2024] for the TSE, using the very same dataset as here. More precisely, the square-root impact law states that
@@
where $\epsilon$ is the sign of the metaorder of total size $Q$, $\Delta p$ is the log mid-price change between just before the first and just after last child order and $\sigma_{D}$ and $V_{D}$ are the contemporaneous daily volatility and exchanged volume. Note again that $\mathcal{I}(Q)$ is independent of the execution time $T$ (see for example [bucci_impact_vol2019] and Fig. 6 below). In the rest of the paper, we will use $({p_{\rm{high}}-p_{\rm{low}}})/{p_{\rm{open}}}$ as a proxy for $\sigma_{D}$.
@@ 3.1 The “double” square-root impact of child orders
Instead of measuring the impact $\mathcal{I}$ of full metaorders, one can measure the partial impact $\mathcal{J}(q,i)$, i.e. the average price difference $\Delta p_{i}$ between just before child order $i+1$ and just before child order $1$, conditional on the size $q$ of child orders and the rank $i$. We find that to a good approximation that impact is proportional to both $\sqrt{q}$ and $\sqrt{i}$:
@@
  • The fit is very good up to $i=50$, beyond which another regime appears to set in, where impact saturates. This however only concerns a small fraction of large metaorders, for which conditioning effects should be taken into account (e.g. large prevailing liquidity at the opposite best, see [patzelt2018universal]). An alternative interpretation is that such large metaorders are detected by the market, triggering the influx of opposing limit orders. Impact saturation for large $Q$ has been reported elsewhere as well, see e.g. [zarinelli2015beyond, bucci2019crossover].
  • When fitting with a more general power-law $(i+i_{0})^{1-\beta}-(i_{0})^{1-\beta}$, the optimal value of $\beta$ is found to be $0.48$, i.e. very close to a square-root. [^2] The exponent $\beta$ is defined as the decay exponent of the propagator, as $|t_{i}-t_{j}|^{-\beta}$, see [bouchaud2003fluctuations, bouchaud2018trades]. Alternatively, imposing $i_{0}=0$ yields $0.71-\beta=$ 0.7, but the fit is clearly worse for small times, see Fig. 4.
  • When $i\gg i_{0}$, one finds that the temporal profile of the impacted price behaves as $\sqrt{i}$, a result already reported in [moro2009market, zarinelli2015beyond, donier2015million].
  • When $i=N$ and using $Q=qN$ one finds
@@
    allowing one to recover exactly the square-root impact law, Eq. (1), up to a weakly varying $N$-dependent factor that increases from $0.31$ to $1$ as $N$ goes from $2$ to $\infty$ (when $i_{0}=4$).

### v2

# The “double” square-root law: Evidence for the mechanical origin of market impact using Tokyo Stock Exchange data

@@ 1 Introduction
The square-root law for price impact is arguably one of the most robust empirical regularities discovered in the last 30 years – see [1, 2, 3, 4, 5, 6] and [7, 8] for reviews. It states that when executing a buy (resp. sell) meta-order of total size $Q$, sliced and diced into $N$ child orders of size $q=Q/N$, the price on average moves up (down) by an amount proportional to $\sqrt{Q}$. Price impact is, quite remarkably, found to be approximately independent of both $N$ and of the total time $T$ needed to achieve full execution [9]. In other words, provided the participation rate is not too large, average price impact only depends – to a first approximation – on the total volume traded $Q$, but not on execution schedule [7].
@@
Such a square-root dependence, and its apparent universality across a wide variety of markets [7], is surprising and non-intuitive. The classical Kyle model would rather predict a linear dependence of impact on $Q$, with a slope usually called “Kyle’s $\lambda$[10]. Square-root impact can be interpreted as an effective $Q$-dependent $\lambda$ that diverges as $Q^{-1/2}$ when $Q\to 0$, as if there was a dearth of liquidity for small volumes [11].
@@
Several theoretical ideas have been put forth in the literature to explain non-linear impact. Some models predict a concave price impact $Q^{\delta}$ with $\delta\leq 1$ related to the power-law tail exponent $\alpha$ of the executed volume [12] or the power-law tail exponent $\gamma$ of the time autocorrelation of the sign of market orders [13, 14]. However, as recently shown by Sato and Kanazawa [15] using ID-resolved data from the Tokyo Stock Exchange (TSE), the predicted relations between $\delta$ and $\alpha$ or $\gamma$ are not borne out by the data: whereas $\alpha$ and $\gamma$ significantly differ between stocks, exponent $\delta$ remains stubbornly anchored around $\delta=1/2$, i.e. the value corresponding to the square-root law. Another possibility is to invoke the “propagator model” which describes the power-law decay of the impact of individual orders [16, 17, 7]. However, in this case all the concavity comes from a time effect, not from a volume effect – a possibility clearly ruled out, see [9] and Fig. 6 below.
@@
Yet another line of thought builds upon the classical Glosten-Milgrom model [18] and assumes that metaorders are issued by traders possessing information about future prices, as is traditional in the economics literature [19, 20]. Market makers try to detect these informed traders, but do not know their actual participation rate in the total order flow. Using Bayesian arguments, Saddier & Marsili [21] then derive a square-root impact law under certain (weak) conditions and predict that the impact will decay after the end of execution as $t^{-1/2}$ when $t\gg T$.
@@
This last prediction is shared with the Locally Linear Order Book (LLOB) model, which is based on a dynamical theory of “latent” liquidity, i.e. volumes that are intended to be exchanged but are not necessarily visible in the order book [22, 7]. Plausible assumptions about the dynamics of such liquidity lead to a generic “V-shape” in the vicinity of the current price, meaning that local liquidity is vanishingly small – in line with our above comment on the divergence of Kyle’s $\lambda$ for small $Q$’s. Incoming buy (sell) trades then face more and more “resistance” as the price moves up (down). The linear increase of liquidity then translates into a square-root behaviour for price impact [22, 11]. However, a more precise formulation of the model suggests the existence of two regimes: a small $Q$ regime where the impact is linear and a larger $Q$ regime where the impact is square root. Whereas such a crossover between the two regimes has indeed been reported in Ref. [23] for very small $Q$’s, the square-root regime sets in for volumes 300 times smaller than predicted by the theory. We will discuss this conundrum more in details below, in the light of our empirical results.
@@
Let us finally mention that the square-root impact law can also be justified using dimensional arguments coupled with additional plausible assumptions, see [24, 25].
@@
Despite its critical importance for both financial microstructure and an asset pricing [26, 27], the very origin of this central phenomenon remains a topic of debate. The universality of the phenomenon suggests a purely mechanical, rather than informational, origin, however this point is controversial and at odds with most of the economics literature on the subject, starting with the famous Kyle model [10]. The aim of the present paper is to give more credence to the “mechanical hypothesis” of the square-root market impact law using an ID-resolved data set from the TSE. Our main results are the following:
@@
  • “Fast” traders, for which the holding period is less than a day, represent between 50% and 60% of the executed market orders. This includes market makers (HFT) and short term traders. Correspondingly, the fraction of market orders executed against “fast” traders represents nearly half of the exchanged volume, a number far too low to vindicate the standard interpretation of the square-root impact within the LLOB framework [28]. Since this empirical fact is inconsistent with the standard LLOB framework, it motivates us to propose a new interpretation of the LLOB framework, as presented in this paper.
  • One can define refill sequences for liquidity providers. One finds that the size of those sequences is also power-law distributed, as was found for metaorders in [29], following the suggestion of Lillo, Mike & Farmer [14].
@@ 2.1 A unique dataset
Our study is based on a dataset from the Tokyo Stock Exchange (TSE), provided by the Japan Exchange Group (JPX) for academic purposes only and already used in [29]. The dataset contains all orders sent to the exchange, with a unique order ID, a virtual server ID, the price and type of the order, the volume and price of the best quotes, for all stocks available on the exchange from 2012 to 2018. Here the virtual server ID is the unit of trading accounts on the TSE. Technically it is not a membership ID (i.e., the corporate level ID) because any trader may have several virtual servers to avoid the submission-number limit during a fixed interval. However, one can reconstruct an effective trader ID, called the Trading Desk, by properly aggregating those virtual server IDs (see [29, 30] for the details). In this paper, the Trading Desks are referred to as trader IDs.
@@
Item (i) is particularly important regarding the claim that impact is an universal mechanism, independent of the type of traders and the information content of the trades [31, 7]. Indeed, access to metaorders data is rare (the Ancerno dataset being one exception [32, 33, 23]), and most of datasets used in the literature are proprietary and may be plagued by conditioning effects [7, 8]. [^1] Note however that CFM data [4] was acquired in such a way to minimize conditioning effects such as those discussed in [7], ch. 12.3, see also [34].
Item (ii) represents an interesting opportunity to categorize traders as market makers (MM), high frequency traders (HFT) or low frequency traders, see 2.3 and understand better their typical impact on the market. For example, the long term debate about the benefits of HFT for market stability has been dramatically improved with this kind of dataset [35, 30]. In addition, these identifiers enabled us to analyze the behavior of liquidity providers, which is only possible if we have access to the ownership of all limit or cancellation orders, see Section 5.
@@ 2.2 General stylized facts about metaorders execution
[…]
The average time between child orders ${\Delta t}$ and the average number of child orders are plotted in Fig. 1 (top graph) as a function of $f=Q/V_{D}$, where $V_{D}$ is the executed volume during the day in shares.. We also show as a shaded region the standard deviation of these quantities. One sees that the average time between child orders mildly increases as a function of $f$, ranging from 252525 secs. for $f=0.01\%$ to 150150150 secs. for $f=1\%$, before saturating or even slightly decreasing for larger values of $f$. Hence the average execution time $T$ increases slightly faster than $Q$ itself, except perhaps for the largest metaorders.
The increase of $\Delta t$ when $Q$ increases is related to the fact that child orders are more and more aggressive in order to complete execution, so traders wisely wait longer before sending the next one, lest they are detected by market makers. This however becomes difficult for large $f$s, because traders are also attempting to execute their metaorders as quickly as possible. Note that translated into total execution time $T$, these results show that for $f=0.01\%$, the typical value of $T$ is 200$\approx$ 200≈ 200 seconds, whereas for large metaorders with $f=10\%$, $T\approx$ 90 minutes. These numbers are however only indicative and the total duration of execution can rise to a full day.
The typical number $N$ of child orders is around 101010 for $f\lesssim 1\%$ before increasing steeply for larger $f$. This reflects the fact that the available volume at the best quotes is relatively small and if traders want to avoid “eating into the book”, then the size $q$ of child orders is also limited, which mechanically pushes the number of child orders up when $Q$ increases. Note that typically the volume available at the best quote is around $10^{-4}V_{D}$ for the most liquid stocks.
@@ 2.3 Time scales & Market ecology
These time scales are relevant to understand order flow and liquidity dynamics, and therefore price impact. Indeed, whereas market makers allow orderly trading by acting as intermediaries between final buyers and final sellers, their inventory constraint prevent them from offering “resistance” to large buy or sell metaorders. Only slow liquidity can counter-act persistent order imbalance [28].
@@
[…]
@@
The conclusion of this study is that while fast traders are dominant in terms of volume, the contribution of slow volumes to trading activity actually of the same order of magnitude. This finding is important since the standard interpretation of the square-root impact law in terms of latent liquidity [22, 28] assumes that slow volumes are a factor 300$\sim$ 300∼ 300 times smaller than $V_{D}$ [23], which is certainly not the case here. In the next Section, we will revisit the empirical evidence for square-root impact and propose a new interpretation of the LLOB model.
@@ 3 Square-root impact: micro-scales & meso-scales
As recalled in the introduction, there is overwhelming empirical evidence for the square-root impact law for metaorders, which has again been confirmed in great detail in [15] for the TSE, using the very same data set as here. More precisely, the square-root impact law states that
@@
where $\epsilon$ is the sign of the metaorder of total size $Q$, $\Delta p$ is the log mid-price change between just before the first and just after last child order and $\sigma_{D}$ and $V_{D}$ are the contemporaneous daily volatility and exchanged volume. Note again that $\mathcal{I}(Q)$ is independent of the execution time $T$ (see for example [9] and Fig. 6 below). In the rest of the paper, we will use $({p_{\rm{high}}-p_{\rm{low}}})/{p_{\rm{open}}}$ as a proxy for $\sigma_{D}$.
@@ 3.1 The “double” square-root impact of child orders
Instead of measuring the impact $\mathcal{I}$ of full metaorders, one can measure the partial impact $\mathcal{J}(q,i)$, i.e. the average price difference $\Delta p_{i}$ between just before child order $i+1$ and just before child order 111, conditional on the size $q$ of child orders and the rank $i$. We find that to a good approximation that impact is proportional to both $\sqrt{q}$ and $\sqrt{i}$:
@@
  • The fit is very good up to $i=50$, beyond which another regime appears to set in, where impact saturates. This however only concerns a small fraction of large metaorders, for which conditioning effects should be taken into account (e.g. large prevailing liquidity at the opposite best, see [38]). An alternative interpretation is that such large metaorders are detected by the market, triggering the influx of opposing limit orders. Impact saturation for large $Q$ has been reported elsewhere as well, see e.g. [32, 23].
  • When fitting with a more general power-law $(i+i_{0})^{1-\beta}-(i_{0})^{1-\beta}$, the optimal value of $\beta$ is found to be 0.480.480.48, i.e. very close to a square-root. [^2] The exponent $\beta$ is defined as the decay exponent of the propagator, as $|t_{i}-t_{j}|^{-\beta}$, see [16, 7]. Alternatively, imposing $i_{0}=0$ yields $1-\beta=0.7$, but the fit is clearly worse for small times, see Fig. 4.
  • When $i\gg i_{0}$, one finds that the temporal profile of the impacted price behaves as $\sqrt{i}$, a result already reported in [39, 32, 5].
  • When $i=N$ and using $Q=qN$ one finds, after injecting $i=N$ in Eq. (2) and factorizing,
@@
    allowing one to recover exactly the square-root impact law, Eq. (1), up to a weakly varying $N$-dependent factor that increases from 0.310.310.31 to 111 as $N$ goes from 222 to $\infty$ (when $i_{0}=4$).